# Zeroth-Order Policy Gradient for Reinforcement Learning from Human Feedback without Reward Inference

### v1

@@ Appendix A Proof Backbone for Theorems: Preference Estimation Error
$\mathbb{E}\left[V_{\mu}(\pi_{{\bm{\theta}}_{0}})\right]-\mathbb{E}\left[V_{\mu}(\pi_{{\bm{\theta}}_{T}})\right]\leq-\frac{\alpha}{8}\sum_{t=0}^{T-1}\mathbb{E}\left[\|\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}})\|_{2}^{2}\right]$
$+\left(\frac{33\alpha d^{2}L^{2}\log M}{\mu^{2}M}+\frac{\alpha\mu^{2}L^{2}d^{2}}{2}+\alpha d^{2}L^{2}\sqrt{\frac{\log M}{M}}+\frac{6\alpha^{2}d^{2}H^{2}L}{\mu^{2}N}\right)T$
@@
$\mu^{2}=\max\left\{9\sqrt{\frac{\log M}{M}},\frac{4H}{L\sqrt{dN}}\right\},\quad\alpha=\frac{1}{12dL}$
then we have the following inequality on the positive drift:
@@
$\mathbb{E}\left[\nabla_{{\bm{\theta}}}V_{\mu}(\pi_{{\bm{\theta}}_{t}})-\left.\hat{{\bm{g}}}_{t}\right|{\mathcal{F}}_{t}\right]$
$=$ $\mathbb{E}\left[\left.\frac{d\left(V(\pi_{{\bm{\theta}}_{t}+\mu{\bm{v}}_{t}})-V(\pi_{{\bm{\theta}}_{t}})\right)}{\mu}{\bm{v}}_{t}\right|{\mathcal{F}}_{t}\right]-\mathbb{E}\left[\left.\frac{d}{\mu}\frac{\sum_{n=1}^{N}\sigma^{-1}(p_{t,n})}{N}{\bm{v}}_{t}\right|{\mathcal{F}}_{t}\right].$
@@
$\mathbb{E}\left[\left.\frac{d}{\mu}\frac{\sum_{n=1}^{N}\sigma^{-1}(p_{t,n})}{N}{\bm{v}}_{t}\right|{\mathcal{F}}_{t}\right]=$ $\frac{d}{\mu}\mathbb{E}\left[\left.\mathbb{E}\left[\left.\frac{\sum_{n=1}^{N}\sigma^{-1}(p_{t,n})}{N}\right|{\bm{v}}_{t}\right]{\bm{v}}_{t}\right|{\mathcal{F}}_{t}\right]$
@@
$=$ $\frac{d}{\mu}\mathbb{E}\left[\left.\mathbb{E}\left[\left.\left(r(\tau_{n,1})-r(\tau_{n,0})\right)\right|{\bm{v}}_{t}\right]{\bm{v}}_{t}\right|{\mathcal{F}}_{t}\right]$
$-\frac{d}{\mu}\mathbb{E}\left[\left.\mathbb{E}\left[\left.\left(\sigma^{-1}(p_{t,n})-[r(\tau_{n,1})-r(\tau_{n,0})]\right)\right|{\bm{v}}_{t}\right]{\bm{v}}_{t}\right|{\mathcal{F}}_{t}\right],$
where the first equality uses law of total expectation. Substitute it back into the bias, we have:
@@
$=$ $V(\pi_{{\bm{\theta}}_{t}+\mu{\bm{v}}_{t}})-\mathbb{E}\left[r(\tau_{n,1})|\tau_{n,1}\sim\pi_{{\bm{\theta}}_{t}+\mu{\bm{v}}_{t}}\right]+\mathbb{E}\left[r(\tau_{n,0})|\tau_{n,0}\sim\pi_{{\bm{\theta}}_{t}}\right]-V(\pi_{{\bm{\theta}}_{t}})$
$=$ $0$
The last equality is due to the definition of the value function and the fact that $\tau_{n,1}$ and $\tau_{n,0}$ are generated by the perturbed policy and the original policy independently. Then, the bias drift only comes from the preference estimation error and we will be able to plug it back and have:
@@
$\leq$ $\frac{2dL}{\mu}\sqrt{\frac{\log M}{M}}\left\|\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}})\right\|_{2}+d^{2}L^{2}\sqrt{\frac{\log M}{M}}$
where the first inequality uses Cauchy-Schwarz inequality and the second inequality uses Lemma. 1. The last inequality is obtained by choosing $M\geq 8(H/L)^{\frac{2}{3}}$.
B.2 Proof of Lemma. 5
@@
Appendix C Proof of Theorem. 2
@@
Then, we analyze the two positive drift terms $\mathsf{Bias}_{t}$ for the first-order gradient bias and $\mathsf{Var}_{t}$ for the gradient noise separately, and then construct a tight upper bound of the Lyapunov drift. The results are summarized in the following two lemmas where the proofs are deferred to Sec. C.1 and Sec. C.2:
@@
$\leq$ $-\frac{\alpha}{2}\|\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}})\|_{2}^{2}+\alpha\mathsf{Bias}_{t}+\frac{L\alpha^{2}}{2}\mathsf{Var}_{t}$
@@
we have the positive first-order drift regarding gradient norm is bounded by:
$\left(\frac{\mu Ld}{2}+\frac{2dL}{\mu}\sqrt{\frac{\log M}{M}}\right)\left\|\nabla_{{\bm{\theta}}}V_{\mu}(\pi_{{\bm{\theta}}_{t}})\right\|_{2}\leq\frac{1}{8}\left\|\nabla_{{\bm{\theta}}}V_{\mu}(\pi_{{\bm{\theta}}_{t}})\right\|_{2},$
@@
$\leq$ $-\frac{\alpha}{8}\|\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}})\|_{2}^{2}+\alpha^{2}\mu^{2}L^{3}d^{2}+\frac{5\alpha^{2}d^{2}L^{3}\log M}{\mu^{2}M}+\frac{6\alpha^{2}d^{2}H^{2}L}{\mu^{2}N}.$
If on the other hand, the gradient is large, then we can upper bound the first-order drift regarding the gradient norm in another way:
@@
$\leq$ $-\frac{\alpha}{8}\sum_{t=0}^{T-1}\mathbb{E}\left[\|\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}})\|_{2}^{2}\right]+\left(5\alpha\mu^{2}L^{2}d^{2}+\frac{65\alpha d^{2}L^{2}\log M}{\mu^{2}M}+\frac{6\alpha^{2}d^{2}H^{2}L}{\mu^{2}N}\right)T$
@@
$\mu^{2}=\max\left\{4\sqrt{\frac{\log M}{M}},\frac{H}{3L\sqrt{dN}}\right\},\quad\alpha=\frac{1}{12dL}$
then we have the following inequality on the positive drift:
@@
$\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}})-\mathbb{E}\left[\left.\hat{{\bm{g}}}_{t}\right|{\mathcal{F}}_{t}\right]$
$=$ $\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}})-\mathbb{E}\left[\left.\mathbb{E}\left[\left.\frac{d}{\mu}\frac{\sum_{n=1}^{N}\sigma^{-1}(p_{t,n})}{N}\right|{\bm{v}}_{t}\right]{\bm{v}}_{t}\right|{\mathcal{F}}_{t}\right]$
@@
where the second inequality uses the fact that conditioned on ${\bm{v}}_{t}$, each trajectory pair $(\tau_{n,1},\tau_{n,0})$ is independently generated from policies $(\pi_{{\bm{\theta}}_{t}+\mu{\bm{v}}_{t}},\pi_{{\bm{\theta}}_{t}})$. The last inequality is due to the fact that conditioned on the policy $(\pi_{{\bm{\theta}}_{t}+\mu{\bm{v}}_{t}},\pi_{{\bm{\theta}}_{t}})$, the expectation of each reward trajectory is the value function of the policy since $(\tau_{n,1},\tau_{n,0})$ is independently generated from policies $(\pi_{{\bm{\theta}}_{t}+\mu{\bm{v}}_{t}},\pi_{{\bm{\theta}}_{t}})$. Therefore, the $\mathsf{Bias}_{t}$ term is written as follows:
@@
$\langle\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}}),\mathsf{Bias}_{t,2}\rangle$
$=$ $\frac{d}{\mu}\mathbb{E}\left[\left.\mathbb{E}\left[\left.\left(\sigma^{-1}(p_{t,n})-[r(\tau_{n,1})-r(\tau_{n,0})]\right)\right|\tau_{n,1},\tau_{n,0}\right]\langle\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}}),{\bm{v}}_{t}\rangle\right|{\mathcal{F}}_{t}\right]$
@@
$\frac{1}{d}\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}})=\mathbb{E}_{{\bm{i}}_{t},\bm{\lambda}_{t}}\left[\langle\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}}),{\bm{v}}_{t}\rangle{\bm{v}}_{t}\right]$
@@
where the first equality is by the definition of ${\bm{v}}_{t}$, the third equality is by law of total expectation. For $j\neq l$, we can use the independence between different coordinates of $\bm{\lambda}_{t}$. Since the expectation of $\bm{\lambda}_{t}$ is zero, we have:
@@
$\langle\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}}),\mathsf{Bias}_{t,1}\rangle$
$=$ $\left\langle\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}}),\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}})-\frac{d}{\mu}\mathbb{E}\left[\left.\left(V(\pi_{{\bm{\theta}}_{t}+\mu{\bm{v}}_{t}})-V(\pi_{{\bm{\theta}}_{t}})\right){\bm{v}}_{t}\right|{\mathcal{F}}_{t}\right]\right\rangle$
@@
$\langle\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}}),\mathsf{Bias}_{t,1}\rangle\leq$ $\frac{d}{\mu}\mathbb{E}\left[\left.\frac{\mu^{2}L}{2}\left\|\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}})\right\|_{2}\right|{\mathcal{F}}_{t}\right]$
$=$ $\frac{\mu Ld}{2}\left\|\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}})\right\|_{2}.$
@@
$\leq$ $\frac{2d^{2}}{\mu^{2}}\mathbb{E}\left[\left.\left\|\left(V(\pi_{{\bm{\theta}}_{t}+\mu{\bm{v}}_{t}})-V(\pi_{{\bm{\theta}}_{t}})-\langle\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}}),\mu{\bm{v}}_{t}\rangle\right){\bm{v}}_{t}\right\|_{2}^{2}\right|{\mathcal{F}}_{t}\right]$
$+\frac{2d^{2}}{\mu^{2}}\mathbb{E}\left[\left.\left\|\langle\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}}),\mu{\bm{v}}_{t}\rangle{\bm{v}}_{t}\right\|_{2}^{2}\right|{\mathcal{F}}_{t}\right].$
@@
Notice that ${\bm{v}}_{t}$ is a vector which on the coordinates chosen by ${\bm{i}}_{t}$, it is either $-1/\sqrt{K}$ or $1/\sqrt{K}$, on other coordinates, it is zero. So ${\bm{v}}_{t}{\bm{v}}_{t}^{\top}$ will result in a diagonal matrix where on position $(i_{t,1},i_{t,1}),\cdots,(i_{t,K},i_{t,K})$ the element is $1/K$ and 00 otherwise. Since each coordinate $i$ will be chosen in each iteration in ${\bm{i}}_{t}$ with probability $K/d$, taking an expectation over ${\bm{v}}_{t}{\bm{v}}_{t}^{\top}$ will result in a diagonal matrix where the diagonal elements are $1/d$. Therefore, we have:

### v2

@@ Appendix B Proof Backbone for Theorems: Preference Estimation Error
The last inequality uses $\alpha\leq(12dL)^{-1}$ and $d\geq 4$. So by taking a maximum of both bounds and taking an expectation, we have:
@@
$\leq$ $-\frac{\alpha}{8}\sum_{t=0}^{T-1}\mathbb{E}\left[\|\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}})\|_{2}^{2}\right]+\left(\frac{33\alpha d^{2}L^{2}\log M}{\mu^{2}M}+\frac{\alpha\mu^{2}L^{2}d^{2}}{2}+\alpha d^{2}L^{2}\sqrt{\frac{\log M}{M}}+\frac{6\alpha^{2}d^{2}H^{2}L}{\mu^{2}N}\right)T.$
@@
$\mu^{2}=\max\left\{9\sqrt{\frac{\log M}{M}},\frac{4H}{L\sqrt{dN}}\right\},\quad\alpha=\frac{1}{12dL}.$
Then, we have the following inequality on the positive drift:
@@
$\mathbb{E}\left[\nabla_{{\bm{\theta}}}V_{\mu}(\pi_{{\bm{\theta}}_{t}})-\left.\hat{{\bm{g}}}_{t}\right|{\mathcal{F}}_{t}\right]=$ $\mathbb{E}\left[\left.\frac{d\left(V(\pi_{{\bm{\theta}}_{t}+\mu{\bm{v}}_{t}})-V(\pi_{{\bm{\theta}}_{t}})\right)}{\mu}{\bm{v}}_{t}\right|{\mathcal{F}}_{t}\right]-\mathbb{E}\left[\left.\frac{d}{\mu}\frac{\sum_{n=1}^{N}\sigma^{-1}(p_{t,n})}{N}{\bm{v}}_{t}\right|{\mathcal{F}}_{t}\right].$
@@
$\mathbb{E}\left[\left.\frac{d}{\mu}\frac{\sum_{n=1}^{N}\sigma^{-1}(p_{t,n})}{N}{\bm{v}}_{t}\right|{\mathcal{F}}_{t}\right]$
$=$ $\frac{d}{\mu}\mathbb{E}\left[\left.\mathbb{E}\left[\left.\frac{\sum_{n=1}^{N}\sigma^{-1}(p_{t,n})}{N}\right|{\bm{v}}_{t}\right]{\bm{v}}_{t}\right|{\mathcal{F}}_{t}\right]$
@@
$=$ $\frac{d}{\mu}\mathbb{E}\left[\left.\mathbb{E}\left[\left.\left(r(\tau_{n,1})-r(\tau_{n,0})\right)\right|{\bm{v}}_{t}\right]{\bm{v}}_{t}\right|{\mathcal{F}}_{t}\right]-\frac{d}{\mu}\mathbb{E}\left[\left.\mathbb{E}\left[\left.\left(\sigma^{-1}(p_{t,n})-[r(\tau_{n,1})-r(\tau_{n,0})]\right)\right|{\bm{v}}_{t}\right]{\bm{v}}_{t}\right|{\mathcal{F}}_{t}\right],$
where the first equality uses the law of total expectation. The second equality uses the independent nature of the trajectories generated by the same policy pair. In the last inequality, we aim to bound the difference between the reward difference estimate from human preference and the ground-truth reward difference, which can be constructed through concentration inequalities. Substitute it back into the bias, we have:
@@
$=$ $V(\pi_{{\bm{\theta}}_{t}+\mu{\bm{v}}_{t}})-\mathbb{E}\left[r(\tau_{n,1})|\tau_{n,1}\sim\pi_{{\bm{\theta}}_{t}+\mu{\bm{v}}_{t}}\right]+\mathbb{E}\left[r(\tau_{n,0})|\tau_{n,0}\sim\pi_{{\bm{\theta}}_{t}}\right]-V(\pi_{{\bm{\theta}}_{t}})=0.$
The last equality is due to the definition of the value function where trajectory returns are unbiased estimates of the value functions and the fact that $\tau_{n,1}$ and $\tau_{n,0}$ are generated by the perturbed policy and the original policy independently. Then, the bias drift only comes from the preference estimation error and we will be able to plug it back and have:
@@
$\leq$ $\frac{2dL}{\mu}\sqrt{\frac{\log M}{M}}\left\|\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}})\right\|_{2}+d^{2}L^{2}\sqrt{\frac{\log M}{M}},$
where the first inequality uses Cauchy-Schwarz inequality and the second inequality uses Lemma. 1. The second last inequality is obtained by choosing $M\geq 8(H/L)^{\frac{2}{3}}$. The final inequality is obtained by applying the third property of Lemma. 3.
C.2 Proof of Lemma. 5
@@
Appendix D Proof of Theorem. 2
@@
Then, we analyze the two positive drift terms $\mathsf{Bias}_{t}$ for the first-order gradient bias and $\mathsf{Var}_{t}$ for the gradient noise separately and then construct a tight upper bound of the Lyapunov drift. The results are summarized in the following two lemmas where the proofs are deferred to Sec. D.1 and Sec. D.2:
@@
$\mathbb{E}\left[V(\pi_{{\bm{\theta}}_{t}})-V(\pi_{{\bm{\theta}}_{t+1}})|{\mathcal{F}}_{t}\right]\leq$ $-\frac{\alpha}{2}\|\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}})\|_{2}^{2}+\alpha\mathsf{Bias}_{t}+\frac{L\alpha^{2}}{2}\mathsf{Var}_{t}$
@@
$\mathbb{E}\left[V(\pi_{{\bm{\theta}}_{t}})-V(\pi_{{\bm{\theta}}_{t+1}})|{\mathcal{F}}_{t}\right]\leq$ $-\frac{\alpha}{2}\|\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}})\|_{2}^{2}+\alpha\left(\frac{\mu Ld}{2}+\frac{2dL}{\mu}\sqrt{\frac{\log M}{M}}\right)\left\|\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}})\right\|_{2}$
@@
we have the positive first-order drift regarding the gradient norm is bounded by:
$\left(\frac{\mu Ld}{2}+\frac{2dL}{\mu}\sqrt{\frac{\log M}{M}}\right)\left\|\nabla_{{\bm{\theta}}}V_{\mu}(\pi_{{\bm{\theta}}_{t}})\right\|_{2}\leq\frac{1}{8}\left\|\nabla_{{\bm{\theta}}}V_{\mu}(\pi_{{\bm{\theta}}_{t}})\right\|_{2}^{2},$
@@
$\mathbb{E}\left[V(\pi_{{\bm{\theta}}_{t}})-V(\pi_{{\bm{\theta}}_{t+1}})|{\mathcal{F}}_{t}\right]\leq$ $-\frac{\alpha}{8}\|\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}})\|_{2}^{2}+\alpha^{2}\mu^{2}L^{3}d^{2}+\frac{5\alpha^{2}d^{2}L^{3}\log M}{\mu^{2}M}+\frac{6\alpha^{2}d^{2}H^{2}L}{\mu^{2}N}.$
If on the other hand, the gradient is small, i.e., when we have:
$\left\|\nabla_{{\bm{\theta}}}V_{\mu}(\pi_{{\bm{\theta}}_{t}})\right\|_{2}\leq 8\left(\frac{\mu Ld}{2}+\frac{2dL}{\mu}\sqrt{\frac{\log M}{M}}\right),$
and then we can upper bound the first-order drift regarding the gradient norm in another way:
@@
$\leq$ $-\frac{\alpha}{8}\sum_{t=0}^{T-1}\mathbb{E}\left[\|\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}})\|_{2}^{2}\right]+\left(5\alpha\mu^{2}L^{2}d^{2}+\frac{65\alpha d^{2}L^{2}\log M}{\mu^{2}M}+\frac{6\alpha^{2}d^{2}H^{2}L}{\mu^{2}N}\right)T.$
@@
$\mu^{2}=\max\left\{4\sqrt{\frac{\log M}{M}},\frac{H}{3L\sqrt{dN}}\right\},\quad\alpha=\frac{1}{12dL}.$
Then, we have the following inequality on the positive drift:
@@
$\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}})-\mathbb{E}\left[\left.\hat{{\bm{g}}}_{t}\right|{\mathcal{F}}_{t}\right]=$ $\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}})-\mathbb{E}\left[\left.\mathbb{E}\left[\left.\frac{d}{\mu}\frac{\sum_{n=1}^{N}\sigma^{-1}(p_{t,n})}{N}\right|{\bm{v}}_{t}\right]{\bm{v}}_{t}\right|{\mathcal{F}}_{t}\right]$
@@
where the second inequality uses the fact that conditioned on ${\bm{v}}_{t}$, each trajectory pair $(\tau_{n,1},\tau_{n,0})$ is independently generated from policies $(\pi_{{\bm{\theta}}_{t}+\mu{\bm{v}}_{t}},\pi_{{\bm{\theta}}_{t}})$. The last inequality is since conditioned on the policy $(\pi_{{\bm{\theta}}_{t}+\mu{\bm{v}}_{t}},\pi_{{\bm{\theta}}_{t}})$, the expectation of each reward trajectory is the value function of the policy since $(\tau_{n,1},\tau_{n,0})$ is independently generated from policies $(\pi_{{\bm{\theta}}_{t}+\mu{\bm{v}}_{t}},\pi_{{\bm{\theta}}_{t}})$. Therefore, the $\mathsf{Bias}_{t}$ term is written as follows:
@@
$\langle\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}}),\mathsf{Bias}_{t,2}\rangle=$ $\frac{d}{\mu}\mathbb{E}\left[\left.\mathbb{E}\left[\left.\left(\sigma^{-1}(p_{t,n})-[r(\tau_{n,1})-r(\tau_{n,0})]\right)\right|\tau_{n,1},\tau_{n,0}\right]\langle\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}}),{\bm{v}}_{t}\rangle\right|{\mathcal{F}}_{t}\right]$
@@
$\frac{1}{d}\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}})=\mathbb{E}_{{\bm{i}}_{t},\bm{\lambda}_{t}}\left[\langle\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}}),{\bm{v}}_{t}\rangle{\bm{v}}_{t}\right],$
@@
where the first equality is by the definition of ${\bm{v}}_{t}$, and the third equality is by the law of total expectation. For $j\neq l$, we can use the independence between different coordinates of $\bm{\lambda}_{t}$. Since the expectation of $\bm{\lambda}_{t}$ is zero, we have:
@@
$\langle\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}}),\mathsf{Bias}_{t,1}\rangle=$ $\left\langle\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}}),\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}})-\frac{d}{\mu}\mathbb{E}\left[\left.\left(V(\pi_{{\bm{\theta}}_{t}+\mu{\bm{v}}_{t}})-V(\pi_{{\bm{\theta}}_{t}})\right){\bm{v}}_{t}\right|{\mathcal{F}}_{t}\right]\right\rangle$
@@
$\langle\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}}),\mathsf{Bias}_{t,1}\rangle\leq$ $\frac{d}{\mu}\mathbb{E}\left[\left.\frac{\mu^{2}L}{2}\left\|\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}})\right\|_{2}\right|{\mathcal{F}}_{t}\right]=\frac{\mu Ld}{2}\left\|\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}})\right\|_{2}.$
@@
$\leq$ $\frac{2d^{2}}{\mu^{2}}\mathbb{E}\left[\left.\left\|\left(V(\pi_{{\bm{\theta}}_{t}+\mu{\bm{v}}_{t}})-V(\pi_{{\bm{\theta}}_{t}})-\langle\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}}),\mu{\bm{v}}_{t}\rangle\right){\bm{v}}_{t}\right\|_{2}^{2}\right|{\mathcal{F}}_{t}\right]+\frac{2d^{2}}{\mu^{2}}\mathbb{E}\left[\left.\left\|\langle\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}}),\mu{\bm{v}}_{t}\rangle{\bm{v}}_{t}\right\|_{2}^{2}\right|{\mathcal{F}}_{t}\right].$
@@
Notice that ${\bm{v}}_{t}$ is a vector such that on the coordinates chosen by ${\bm{i}}_{t}$, the element is either $-1/\sqrt{K}$ or $1/\sqrt{K}$, and on other coordinates, it is zero. So ${\bm{v}}_{t}{\bm{v}}_{t}^{\top}$ will result in a diagonal matrix where on position $(i_{t,1},i_{t,1}),\cdots,(i_{t,K},i_{t,K})$ the element is $1/K$ and 00 otherwise. Since each coordinate $i$ will be chosen in each iteration in ${\bm{i}}_{t}$ with probability $K/d$, taking an expectation over ${\bm{v}}_{t}{\bm{v}}_{t}^{\top}$ will result in a diagonal matrix where the diagonal elements are $1/d$. Therefore, we have:
@@
So combining three terms, the variance drift can be bounded as: